OpenStudy (anonymous):

Solve on the interval : [0,2pi): 2cscx-3=-5

OpenStudy (dumbcow):

isolate csc x using basic algebra \[\csc x = -1\] now csc is the reciprocal of sin, or 1/sin \[\frac{1}{\sin x} = -1\] isolating sin x \[\sin x = -1\] from the unit circle you find that sin x = -1 at an angle of 3pi/2 \[x = \frac{3 \pi}{2}\] http://www.google.com/imgres?imgurl=http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/720px-Unit_circle_angles_color.svg.png&imgrefurl=http://en.wikipedia.org/wiki/Unit_circle&h=225&w=225&tbnid=NHFRVspjd-vZrM:&zoom=1&tbnh=186&tbnw=186&usg=__sOrTVeIoKl48_TodEan8dIx7UGk=&docid=D09vYD7H2hKF-M&itg=1&sa=X&ei=Og5DU8qjE-ewyQGwr4GQCQ&ved=0CKMBEPwdMAo